# Bootstrapping Vision-language Models for Self-supervised Remote Physiological Measurement

### v1

@@ 4.2 Implementation Details
Hyper-parameters. For facial videos in four datasets, we randomly sample consecutive $F=224$ frames from each video to build the spatio-temporal maps (STMaps) $M^{p}=\left\{m^{p}\right\}$ and $M^{n}=\left\{m^{n}\right\}$ for positive and negative samples, respectively. The dimension of the STMap is $224\times 224\times 3$, ensuring compatibility with the size requirement for the ViT-based vision encoder. The temperature $\tau_{1}$ in equation (2) is set as 0.07 following [25], and $\tau_{2}$ in equation (3) is set as 0.08 following [21]. The masking ratio $b=60\%$ is equivalent to that in [58]. We freeze the weights of the pre-trained LFA module.
@@ 4.3 Evaluation Protocol
TABLE II: Comparison to state of the art on RF and HRV estimation. The results are reported on the UBFC-rPPG. $\uparrow$ indicates that the larger the value is the better it is and $\downarrow$ vice versa. The best supervised approach is marked in shadow, while the best self-supervised approach is marked in bold.
Method UBFC-rPPG PPG RF HRV: LF HRV: HF HRV: LF/HF
@@
POS [12] - 0.109 0.107 0.087 0.171 0.169 0.479 0.171 0.169 0.479 0.405 0.399 0.518
CHROM [14] - 0.086 0.089 0.102 0.243 0.240 0.159 0.243 0.240 0.159 0.655 0.645 0.266
Green [3] - 0.087 0.086 0.111 0.186 0.186 0.280 0.186 0.186 0.280 0.361 0.365 0.492
CVD [19] 0.017 0.018 0.252 0.053 0.065 0.740 0.053 0.065 0.740 0.169 0.168 0.812
rPPGNet [17] 0.030 0.034 0.233 0.071 0.070 0.686 0.071 0.070 0.686 0.212 0.208 0.744
Dual-Gan [5] 0.010 0.010 0.395 0.034 0.035 0.891 0.034 0.035 0.891 0.131 0.136 0.881
Physformer [6] 0.009 0.009 0.413 0.030 0.032 0.895 0.030 0.032 0.895 0.126 0.130 0.893
Gideon et al. [22] 0.061 0.098 0.103 0.091 0.139 0.694 0.091 0.139 0.694 0.525 0.691 0.684
Yue et al. [21] 0.023 0.028 0.351 0.047 0.062 0.769 0.047 0.062 0.769 0.160 0.164 0.831
Contrast-Phys+ [23] 0.085 0.083 0.347 0.096 0.098 0.798 0.096 0.098 0.798 0.391 0.395 0.782
4.4 Results
4.4.1 HR evaluation
We first conduct intra-dataset HR estimation on four datasets. As shown in Table I, we compare the proposed VL-phys with three traditional methods (POS [12], CHROM [14] and Green [3]); eight DNN-based supervised methods (SynRhythm [48], Meta-rppg [49], PulseGan [50], Dual-Gan [5], Physformer [6], Du et al. [7], Dual-TL [51], and ND-DeeprPPG [52]); four DNN-based self-supervised methods (Gideon et al. [22], Contrast-Phys+ [23], Yue et al. [21], and SiNC [53]).
@@ 4.4.1 HR evaluation
TABLE III: Comparison to state of the art on cross-dataset HR estimation. $\uparrow$ indicates that the larger the value is the better it is and $\downarrow$ vice versa. The best supervised approach is marked in shadow, while the best self-supervised approach is marked in bold.
@@
Meta-rppg [49] 6.48 7.97 0.52 5.69 7.74 0.84 6.11 7.58 0.66 4.00 5.98 0.92
PulseGan [50] 2.33 3.62 0.97 4.40 6.35 0.89 2.30 3.50 0.97 3.36 5.11 0.95
Dual-Gan [5] 2.00 3.13 0.97 3.51 4.99 0.93 2.03 3.01 0.97 1.81 2.97 0.99
Physformer [6] 1.97 3.08 0.97 3.46 4.96 0.93 1.93 3.02 0.97 1.99 3.28 0.99
ND-DeeprPPG [52] - - - - - - 0.34 0.98 0.99 0.17 0.35 0.99
Gideon et al. [22] 2.45 3.71 0.93 3.84 5.31 0.91 2.37 3.51 0.95 2.95 4.60 0.97
Yue et al. [21] 2.24 3.40 0.97 3.50 4.95 0.93 2.18 3.20 0.97 2.14 3.37 0.98
Contrast-Phys+ [23] 2.18 3.42 0.97 3.55 5.03 0.92 2.07 3.10 0.97 2.03 3.35 0.99
@@
4.4.2 RF and HRV evaluation
We further conduct intra-dataset RF and HRV estimation on UBFC-rPPG. We compare our approach with state of the art, and present the results in Table II. LF, HF, and LF/HF are three attributes for HRV estimation. We can observe that the proposed method significantly outperforms traditional methods and most DNN-based methods. These findings underscore the capability of our method to yield high-quality rPPG signals with accurate systolic peaks, resulting into superior HRV estimations.
@@ 5.1 Cross-dataset HR evaluation
TABLE IV: HR estimation of VL-phys on four datasets under a semi-supervised setting.
@@
5.2 Semi-supervised HR Evaluation
We then evaluate the performance of VL-phys on four datasets under a semi-supervised setting, as shown in Table IV. VL-phys(90%+10%) indicates that 90% samples in the training set are utilized for self-supervised learning, while the remaining 10% are used for supervised learning. The two types of data are passed through the same backbone, except that the latter also leverages the ground truth supervision into the training. We use the Pearson correlation coefficient-based loss [24] to constrain the same rhythm periodicity between the estimated rPPG and the ground truth. The other variants in Table IV are trained in the same setup. First, we can observe that, in this semi-supervised setting, when a partial of ground truth signals is available, VL-phys shows a significant performance improvement, even surpassing the best-performing supervised methods presented in Table I. For instance, with only 30% of the samples provided with ground truth signals, VL-phys achieves an MAE of 0.13, 0.16, 4.86 and 1.93 across four datasets, respectively, significantly outperforming the best supervised methods. Second, an increase in labeled data from VL-phys to VL-phys(70%+30%) yields superior outcomes, suggesting that more ground truth can effectively enhance the learning of the periodicity of rPPG signals. Moreover, these results also demonstrate that our method can be effectively adapted to a semi-supervised framework and exhibits superior performance in scenarios where some labeled videos are available.
@@ 6 Ablation study
TABLE V: Ablation study on the text modality.
@@
6.1 Text modality
In our method, we for the first time leverage text prompts to enhance the vision encoder’s understanding of frequency-related visual attributes for rPPG estimation. We now investigate the effectiveness of leveraging the text modality. In Table V, we denote VL-phys w/o text as a variant where the text modality is removed from the framework. In this way, the STMaps for positive and negative samples are directly fed into the vision encoder for rPPG estimation without constructing the vision-text pairs. The proposed TVR module, vision-text contrastive loss $L_{vtc}$, and reconstruction loss $L_{r}$ are removed due to the absence of the text modality. The results show a significant increase in MAE, rising to 0.57 on the UBFC-rPPG dataset and 2.88 on the MMVS dataset, attesting the usage of the text modality in this task.
Text-guided visual map reconstruction task. Text modality is leveraged into a generative learning task, i.e. , text-guided visual map reconstruction task, to optimize the vision encoder to reconstruct the C-STMaps. Specifically, a TVR module and a reconstruction loss $L_{r}$ are designed to recover the masked patches of C-STMaps based on the linguistic clues provided by their associated text prompts. If we remove this task from our framework, the MAE and RMSE will be increased for HR estimation (see Table V: VL-phys w/o TVR). This observation suggests that this task can effectively encourage the interaction between vision and text modalities, enabling the vision encoder to capture accurate periodic patterns in temporal variations of skin colors.
Vision-text contrastive loss. We further validate the effectiveness of the vision-text contrastive loss $L_{vtc}$ by ablating it in our framework, see Table V. A clear performance drop can be observed by VL-phys w/o $L_{vtc}$. This justifies our primary motivation of performing vision-text contrastive learning to boost the performance.
TABLE VI: Ablation study on the vision and text encoders.
@@ 6.1 Text modality
VL-phys(CLIP) [25] 0.49 0.79 0.99 2.83 4.07 0.94
VL-phys(COCA) [61] 0.47 0.74 0.99 2.65 3.90 0.95
VL-phys(ALBEF) [62] 0.48 0.80 0.99 2.69 4.01 0.95
VL-phys(BLIP) [39] 0.41 0.71 0.99 2.45 3.81 0.95
VL-phys(BLIP2) [63] 0.41 0.70 0.99 2.47 3.82 0.95
VL-phys(All-in-one) [27] 0.25 0.61 0.99 2.36 3.79 0.95
VL-phys(CLIP-ViP) [64] 0.26 0.65 0.99 2.34 3.78 0.95
@@ 6.2 Vision and text encoders
We employ the pre-trained vision and text encoders from VindLU [26] to extract features from the generated vision-text pairs. Table VI details the performance of variants using encoders from other widely-used VLMs, i.e. , CLIP [25], COCA [61], ALBEF [62], BLIP [39], BLIP2 [63], All-in-one [27], CLIP-ViP [64]. For example, VL-phys(CLIP) indicates a variant in which we use the pre-trained vision and text encoders from CLIP.
First, it is evident that all of these variants outperform the state of the art self-supervised rPPG estimation methods in Table I. This essentially validates that the VLMs fine-tuned with our VL-phy possess more powerful ability to capture the periodic variation of skin color for rPPG estimation.
@@
Pre-trained model weights. We present another variant that does not use the pre-trained weights from VindLU, but instead to initialize the weights of the encoders randomly. This variant is denoted as VL-phys w/o pw in Table VI. The results get worse to 0.55 and 2.39 in MAE on two datasets, respectively. This indicates that pre-training is important for the understanding of the text prompts describing variational visual details.
@@
Textual embedding extraction. We verify the benefits of using text encoder BERT in our framework. Given that BERT has approximately 110 million parameters, we introduce an alternative light-weight variant where the text encoder is only an MLP with five linear layers. This encoder exclusively processes the [ratio] token from the text, and its output is directly employed for vision-text contrastive learning and text-guided visual map reconstruction. This variant is denoted as VL-phys-MLP in Table VI. We can observe severe performance degradation by this variant, attesting the importance of using an established text encoder. Moreover, compared to using the single token of [ratio], our designed text prompt provides richer semantics describing the skin color temporal variation.
@@
Image-guided masked language modeling task. Image-guided masked language modeling task is another kind of generative learning task that could potentially be integrated into our VL-phys. Given a vision-text pair, it aims to predict several randomly masked tokens in the text prompt conditioned on visual tokens and the remaining unmasked text tokens. This task differs from our designed text-guided visual map reconstruction task, which focuses on recovering visual patches. We have incorporated this auxiliary task into our framework to evaluate it. This variant is denoted as VL-phys-MLM in Table VI. It should be noted that popular VLMs typically leverage a multimodal fusion encoder to fuse multimodal cues and reconstruct the masked tokens based on vision-grounded textual embeddings. We employ the pre-trained multimodal fusion encoder from VindLU for this purpose. We can observe a degraded performance for this variant. Similar findings are also noted from VL-phys(BLIP) to VL-phys(BLIP)-MLM. The issue primarily arises due to the uninformative tokens within the text prompts, such as ”the”, ”of”, ”is” and ”as”. When these tokens are masked instead of more informative ones like the [ratio] token, the model tends to predict them using solely linguistic cues, rather than trying to find the answer from the visual modality.
@@
Furthermore, we have improved VL-phys-MLM by setting the masked tokens in the text prompts consistently to the most important [ratio] token, denoted as VL-phys-MLM(r) in Table VI. This variant outperforms VL-phys-MLM and indeed shows some negligible improvement to VL-phys. However, considering the additional computation incurred by the multimodal fusion encoder, we recommend not to use this variant by default.
TABLE VII: Ablation study on loss functions.
@@
TABLE VIII: Ablation study on different text prompts.

### v2

@@ 4.2 Implementation Details
Hyper-parameters. For facial videos in four datasets, we randomly sample consecutive $F=224$ frames from each video to build the spatio-temporal maps (STMaps) $M^{p}=\left\{m^{p}\right\}$ and $M^{n}=\left\{m^{n}\right\}$ for positive and negative samples, respectively. We follow [36] to set the number of detected ROIs in facial videos as $A=25$. The dimension of the STMap is $224\times 224\times 3$, ensuring compatibility with the size requirement for the ViT-based vision encoder. The temperature $\tau_{1}$ in equation (2) is set as 0.07 following [42], and $\tau_{2}$ in equation (3) is set as 0.08 following [71]. The masking ratio $b=60\%$ is equivalent to that in [17]. We freeze the weights of the pre-trained LFA module.
@@ 4.3 Evaluation Protocol
Table 1: Comparison to state of the art on HR estimation. The results are reported on UBFC-rPPG, PURE, VIPL-HR, and MMVS datasets. $\uparrow$ indicates that the larger the value is the better it is and $\downarrow$ vice versa. The best supervised approach is marked in shadow, while the best self-supervised approach is marked in bold.
Method PPG UBFC-rPPG PURE VIPL-HR MMVS
@@
POS - 8.35 10.00 0.24 3.14 10.57 0.95 11.53 17.26 0.30 6.77 9.40 0.82
CHROM - 8.20 9.92 0.27 3.82 6.80 0.97 11.42 16.92 0.28 6.85 9.37 0.82
Green - 6.01 7.87 0.29 4.39 11.60 0.90 - - - 7.13 9.46 0.80
SynRhythm 5.59 6.82 0.72 2.71 4.86 0.98 - - - 4.48 6.52 0.89
Meta-rppg 5.97 7.42 0.53 2.52 4.63 0.98 - - - 4.30 6.20 0.91
PulseGan 1.19 2.10 0.98 2.28 4.29 0.99 - - - 3.52 5.09 0.93
Dual-Gan 0.44 0.67 0.99 0.82 1.31 0.99 4.93 7.68 0.81 3.00 4.27 0.94
Physformer 0.40 0.71 0.99 1.10 1.75 0.99 4.97 7.79 0.78 3.28 4.50 0.93
Du et al. 0.16 0.57 0.99 - - - - - - - - -
Dual-TL 0.17 0.41 0.99 0.37 0.68 0.99 6.38 6.92 0.69 - - -
ND-DeeprPPG 0.31 0.98 0.99 0.18 0.41 0.99 - - - - - -
Gideon et al. 1.85 4.28 0.93 2.32 2.97 0.99 9.80 15.48 0.38 3.43 4.74 0.93
Yue et al. 0.58 0.94 0.99 1.23 2.01 0.99 - - - 2.93 4.16 0.94
Contrast-Phys+ 0.64 1.00 0.99 1.00 1.40 0.99 7.49 14.40 0.49 2.96 4.11 0.94
SiNC 0.59 1.83 0.99 0.61 1.84 0.99 - - - 3.15 4.40 0.93
VL-phys 0.28 0.60 0.99 0.52 0.95 0.99 6.04 8.78 0.71 2.33 3.76 0.95
@@
4.4 Results
4.4.1 HR evaluation
We first conduct intra-dataset HR estimation on four datasets. As shown in Table 1, we compare the proposed VL-phys with three traditional methods (POS [57], CHROM [7] and Green [52]); eight DNN-based supervised methods (SynRhythm [35], Meta-rppg [20], PulseGan [46], Dual-Gan [32], Physformer [67], Du et al. [9], Dual-TL [40], and ND-DeeprPPG [27]); four DNN-based self-supervised methods (Gideon et al. [10], Contrast-Phys+ [49], Yue et al. [71], and SiNC [47]).
{Moreover, it is worth noting that the general performance of comparable methods on the VIPL-HR dataset is clearly inferior to that on the rest datasets. This performance discrepancy arises because the facial videos in the VIPL-HR dataset were collected in less-constrained scenarios, including diverse ambient lighting conditions, and varying camera angles. These challenging factors introduce noises and hinder the capturing of physiological clues for accurate rPPG estimation. In contrast, other datasets were collected under controlled conditions, with consistent lighting and stable camera positioning.
@@ 4.4.1 HR evaluation
4.4.2 RF and HRV evaluation
We further conduct intra-dataset RF and HRV estimation on UBFC-rPPG. We compare our approach with state of the art, and present the results in Table 2. LF, HF, and LF/HF are three attributes for HRV estimation. We can observe that the proposed method significantly outperforms traditional methods and most DNN-based methods. These findings underscore the capability of our method to yield high-quality rPPG signals with accurate systolic peaks, resulting into superior HRV estimations.
Table 2: Comparison to state of the art on RF and HRV estimation. The results are reported on the UBFC-rPPG. $\uparrow$ indicates that the larger the value is the better it is and $\downarrow$ vice versa. The best supervised approach is marked in shadow, while the best self-supervised approach is marked in bold.
Method PPG RF HRV: LF HRV: HF HRV: LF/HF
[…]
POS - 0.109 0.107 0.087 0.171 0.169 0.479 0.171 0.169 0.479 0.405 0.399 0.518
CHROM - 0.086 0.089 0.102 0.243 0.240 0.159 0.243 0.240 0.159 0.655 0.645 0.266
Green - 0.087 0.086 0.111 0.186 0.186 0.280 0.186 0.186 0.280 0.361 0.365 0.492
CVD 0.017 0.018 0.252 0.053 0.065 0.740 0.053 0.065 0.740 0.169 0.168 0.812
rPPGNet 0.030 0.034 0.233 0.071 0.070 0.686 0.071 0.070 0.686 0.212 0.208 0.744
Dual-Gan 0.010 0.010 0.395 0.034 0.035 0.891 0.034 0.035 0.891 0.131 0.136 0.881
Physformer 0.009 0.009 0.413 0.030 0.032 0.895 0.030 0.032 0.895 0.126 0.130 0.893
Gideon et al. 0.061 0.098 0.103 0.091 0.139 0.694 0.091 0.139 0.694 0.525 0.691 0.684
Yue et al. 0.023 0.028 0.351 0.047 0.062 0.769 0.047 0.062 0.769 0.160 0.164 0.831
Contrast-Phys+ 0.085 0.083 0.347 0.096 0.098 0.798 0.096 0.098 0.798 0.391 0.395 0.782
Table 3: Comparison to state of the art on cross-dataset HR estimation. $\uparrow$ indicates that the larger the value is the better it is and $\downarrow$ vice versa. The best supervised approach is marked in shadow, while the best self-supervised approach is marked in bold.
@@ 4.4.2 RF and HRV evaluation
Meta-rppg 6.48 7.97 0.52 5.69 7.74 0.84 6.11 7.58 0.66 4.00 5.98 0.92
PulseGan 2.33 3.62 0.97 4.40 6.35 0.89 2.30 3.50 0.97 3.36 5.11 0.95
Dual-Gan 2.00 3.13 0.97 3.51 4.99 0.93 2.03 3.01 0.97 1.81 2.97 0.99
Physformer 1.97 3.08 0.97 3.46 4.96 0.93 1.93 3.02 0.97 1.99 3.28 0.99
ND-DeeprPPG - - - - - - 0.34 0.98 0.99 0.17 0.35 0.99
Gideon et al. 2.45 3.71 0.93 3.84 5.31 0.91 2.37 3.51 0.95 2.95 4.60 0.97
Yue et al. 2.24 3.40 0.97 3.50 4.95 0.93 2.18 3.20 0.97 2.14 3.37 0.98
Contrast-Phys+ 2.18 3.42 0.97 3.55 5.03 0.92 2.07 3.10 0.97 2.03 3.35 0.99
@@ 5.1 Cross-dataset HR evaluation
Table 4: HR estimation of VL-phys on four datasets under semi-supervised and fully-supervised settings.
@@
VL-phys(60%+40%) 0.12 0.30 0.99 0.16 0.44 0.99 4.86 7.55 0.82 1.93 2.96 0.97
VL-phys(50%+50%) 0.12 0.28 0.99 0.15 0.42 0.99 4.85 7.54 0.82 1.92 2.95 0.97
VL-phys(0%+100%) 0.11 0.27 0.99 0.13 0.40 0.99 4.83 7.51 0.83 1.88 2.91 0.97
5.2 Semi-supervised and fully-supervised HR Evaluation
We then evaluate the performance of VL-phys on four datasets under the semi-supervised setting and the fully-supervised setting, as shown in Table 4. For instance, VL-phys (90%+10%) indicates that 90% samples in the training set are utilized for self-supervised learning, while the remaining 10% are used for supervised learning. The two types of data are passed through the same backbone, except that the latter also leverages the ground truth supervision into the training. We use the Pearson correlation coefficient-based loss [28] to constrain the same rhythm periodicity between the estimated rPPG and the ground truth. VL-phys (0%+100%) indicates the fully-supervised setting, in which all samples in the training set are leveraged for supervised learning. First, we can observe that, in the semi-supervised setting, when a partial of ground truth signals is available, VL-phys shows a significant performance improvement, even surpassing the best-performing supervised methods presented in Table 1. For instance, with only 30% of the samples provided with ground truth signals, VL-phys achieves an MAE of 0.13, 0.16, 4.86 and 1.93 across four datasets, respectively, significantly outperforming the best supervised methods. Second, an increase in labeled data from VL-phys to VL-phys (50%+50%) yields superior outcomes, suggesting that more ground truth can effectively enhance the learning of the periodicity of rPPG signals. Third, the fully-supervised variant VL-phys (0%+100%) achieves the best performance. These results demonstrate that our method can be effectively adapted to a semi-supervised or fully-supervised framework and exhibits superior performance in scenarios where some labeled videos are available.
Table 5: Quantitative performance on subjects with or without makeup.
Method MMVS-w/ makeup MMVS-w/o makeup
[…]
Gideon et al. 3.90 5.62 0.92 3.39 4.69 0.93
VL-phys 2.47 3.78 0.94 2.15 3.30 0.95
Refer to caption Refer to caption
(a) (b)
Figure 7: Qualitative performance on subjects with makeup.
5.3 Effect of Makeup
We further assess the impact of makeup on our method. Following [71], we divide the MMVS dataset into two subsets: MMVS-w/ makeup and MMVS-w/o makeup. Each subset comprises a training set and a test set, which are used for model training and evaluation, respectively. The HR estimation results are summarized in Table 5. We observe that the performance on MMVS-w/o makeup surpasses that on MMVS-w/ makeup: e.g. the MAE is 2.15 for the former and 2.47 for the latter. Additionally, compared to [10], which exhibits a MAE difference of 0.51 between the two subsets, the influence of makeup on our method is smaller (i.e. 0.32 MAE difference). Fig. 7 shows two estimated rPPG signals and their ground truth signals from subjects with makeup. We can see that some signal peaks are not generated (red arrows). This is because makeup can cover the skin color temporal variations, thereby hindering the extraction of physiological clues for accurate rPPG estimation.
@@ 6 Ablation study
6.1 Text modality
In our method, we for the first time leverage text prompts to enhance the vision encoder’s understanding of frequency-related visual attributes for rPPG estimation. We now investigate the effectiveness of leveraging the text modality. In Table 6, we denote VL-phys w/o text as a variant where the text modality is removed from the framework. In this way, the STMaps for positive and negative samples are directly fed into the vision encoder for rPPG estimation without constructing the vision-text pairs. The proposed TVR module, vision-text contrastive loss $L_{vtc}$, and reconstruction loss $L_{r}$ are removed due to the absence of the text modality. The results show a significant increase in MAE, rising to 0.57 on the UBFC-rPPG dataset and 2.88 on the MMVS dataset, attesting the usage of the text modality in this task.
Text-guided visual reconstruction task. Text modality is leveraged into the frequency-related generative learning task, i.e. , text-guided visual reconstruction task, to optimize the vision encoder to reconstruct the C-STMaps. Specifically, a TVR module and a reconstruction loss $L_{r}$ are designed to recover the masked patches of C-STMaps based on the linguistic clues provided by their associated text prompts. If we remove this task from our framework, the MAE and RMSE will be increased for HR estimation (see Table 6: VL-phys w/o TVR). This observation suggests that this task can effectively encourage the interaction between vision and text modalities, enabling the vision encoder to capture accurate periodic patterns in temporal variations of skin colors.
Vision-text contrastive loss. We further validate the effectiveness of the vision-text contrastive loss $L_{vtc}$ by ablating it in our framework, see Table 6. A clear performance drop can be observed by VL-phys w/o $L_{vtc}$. This justifies our primary motivation of performing vision-text contrastive learning to boost the performance.
Table 6: Ablation study on the text modality.
@@ 6.1 Text modality
Table 7: Ablation study on the vision and text encoders.
@@
VL-phys(CLIP) 0.49 0.79 0.99 2.83 4.07 0.94
VL-phys(COCA) 0.47 0.74 0.99 2.65 3.90 0.95
VL-phys(ALBEF) 0.48 0.80 0.99 2.69 4.01 0.95
VL-phys(BLIP) 0.41 0.71 0.99 2.45 3.81 0.95
VL-phys(BLIP2) 0.41 0.70 0.99 2.47 3.82 0.95
VL-phys(All-in-one) 0.25 0.61 0.99 2.36 3.79 0.95
VL-phys(CLIP-ViP) 0.26 0.65 0.99 2.34 3.78 0.95
@@ 6.2 Vision and text encoders
We employ the pre-trained vision and text encoders from VindLU [6] to extract features from the generated vision-text pairs. Table 7 details the performance of variants using encoders from other widely-used VLMs, i.e. , CLIP [42], COCA [64], ALBEF [21], BLIP [22], BLIP2 [23], All-in-one [55], CLIP-ViP [59]. For example, VL-phys(CLIP) indicates a variant in which we use the pre-trained vision and text encoders from CLIP.
First, it is evident that all of these variants outperform the state of the art self-supervised rPPG estimation methods in Table 1. This essentially validates that the VLMs fine-tuned with our VL-phy possess more powerful ability to capture the periodic variation of skin color for rPPG estimation.
@@
Pre-trained model weights. We present another variant that does not use the pre-trained weights from VindLU, but instead to initialize the weights of the encoders randomly. This variant is denoted as VL-phys w/o pw in Table 7. The results get worse to 0.55 and 2.39 in MAE on two datasets, respectively. This indicates that pre-training is important for the understanding of the text prompts describing variational visual details.
@@
Textual embedding extraction. We verify the benefits of using text encoder BERT in our framework. Given that BERT has approximately 110 million parameters, we introduce an alternative light-weight variant where the text encoder is only an MLP with five linear layers. This encoder exclusively processes the [ratio] token from the text, and its output is directly employed for vision-text contrastive learning and text-guided visual reconstruction. This variant is denoted as VL-phys-MLP in Table 7. We can observe severe performance degradation by this variant, attesting the importance of using an established text encoder. Moreover, compared to using the single token of [ratio], our designed text prompt provides richer semantics describing the skin color temporal variation.
@@
Image-guided masked language modeling task. Image-guided masked language modeling task is another kind of generative learning task that could potentially be integrated into our VL-phys. Given a vision-text pair, it aims to predict several randomly masked tokens in the text prompt conditioned on visual tokens and the remaining unmasked text tokens. This task differs from our designed text-guided visual reconstruction task, which focuses on recovering visual patches. We have incorporated this auxiliary task into our framework to evaluate it. This variant is denoted as VL-phys-MLM in Table 7. It should be noted that popular VLMs typically leverage a multimodal fusion encoder to fuse multimodal cues and reconstruct the masked tokens based on vision-grounded textual embeddings. We employ the pre-trained multimodal fusion encoder from VindLU for this purpose. We can observe a degraded performance for this variant. Similar findings are also noted from VL-phys(BLIP) to VL-phys(BLIP)-MLM. The issue primarily arises due to the uninformative tokens within the text prompts, such as ”the”, ”of”, ”is” and ”as”. When these tokens are masked instead of more informative ones like the [ratio] token, the model tends to predict them using solely linguistic cues, rather than trying to find the answer from the visual modality.
@@
Furthermore, we have improved VL-phys-MLM by setting the masked tokens in the text prompts consistently to the most important [ratio] token, denoted as VL-phys-MLM(r) in Table 7. This variant outperforms VL-phys-MLM and indeed shows some negligible improvement to VL-phys. However, considering the additional computation incurred by the multimodal fusion encoder, we recommend not to use this variant by default.
Fine-tuning methods. During the training stage of VL-phys, we fine-tune the pre-trained vision and text encoders of VLM using the full fine-tuning strategy. We present three variants that do not use the full fine-tuning strategy, but instead use parameter-efficient fine-tuning techniques for network optimization, including prompt-tuning (i.e. VPT [13] and CoOp [74]) and adapter-tuning [4]. These variants are denoted as VL-phys (VPT), VL-phys (CoOp), and VL-phys (Adapter) in Table 7. We can see they achieve inferior performance compared to the original VL-phys. This can be attributed to the difficulty of fine-tuning the encoders to effectively understand frequency-related attributes. Unlike parameter-efficient fine-tuning, full fine-tuning enables the encoders to comprehensively digest the frequency-related knowledge of skin color temporal variation.
Table 8: Ablation study on loss functions.
@@
6.3 Losses
Frequency contrastive loss. Frequency contrastive loss $L_{fc}$ is the most important loss in existing self-supervised remote physiological measurement methods [62, 71, 10, 49]. It enforces the signal frequency similarities among positive samples and dissimilarities between positive and negative samples in the feature space. In Table 8 we use VL-phys w/o $L_{fc}$ to denote a variant without using $L_{fc}$ in our framework. We can see that the MAE and RMSE are substantially increased while $\rho$ is substantially decreased from the original framework, e.g. , +2.42, +2.83 and -0.09 on MMVS, which proves the benefits of adding $L_{fc}$.
Frequency ranking loss. We verify the effectiveness of the proposed frequency ranking loss $L_{fr}$ (Sec. 3.6). First, as shown in Table 8, omitting $L_{fr}$ leads to deteriorated results on both datasets. Next, $L_{fr}$ optimizes the relative frequency ranks of predicted positive and negative signals rather than their absolute frequency differences to avoid overfitting. If we instead constrains the signal frequency ratios between these signals to be consistent with the absolute ground truth ratios $R$, denoted by $L_{fr}\rightarrow L_{afr}$ in Table 8, the performance gets worse, e.g. , +0.05 and +0.09 in MAE on two datasets, respectively.
Table 9: Ablation study on different text prompts.